# On Partition Functions for Time-Inhomogeneous Branching Random Walks

### 15 Mar 2026 v1

@@ Abstract
We establish the phase transition and universality for the partition function of time-inhomogeneous branching random walks (BRWs) with decreasing variance increment, a model related to two-dimensional directed polymers. By modifying Berestycki’s GMC framework (removing martingale property dependency) and adapting it to discrete BRWs, we prove that the critical value of time-inhomogeneous BRWs coincides with that of time-homogeneous ones, and the partition functions converge in $L^{1}$ in the subcritical regime. We also extend the universality at the critical parameter, showing the same decay rate of partition functions. Our approach reveals the potential for such framework in GMC, which provides a new technical path for martingale-free processes and random fields beyond log-correlated. Finally, we raise some open problems related to GMC beyond log-correlations, branching Brownian motions and directed polymers.
@@ 1 Introduction
Branching random walks (BRWs) are a natural model in probability theory, motivated by their biological applications, that describe the evolution of a population of particles with spatial motion. Much is known about the phase transition of their partition function (the so-called Biggins martingale); see [17]. More recently, their time-inhomogeneous counterpart, first introduced in [14], initiates the study of the asymptotic behaviors of BRWs with increments that exhibit a macroscopic and monotonic change in variance, revealing differences in extreme value statistics. The series of works on this model raises a more fundamental question: How do environmental fluctuations impact the asymptotic behavior of BRWs? In the present paper, we attempt to answer this question from a qualitative perspective by studying its partition function.
@@
The partition function of time-inhomogeneous BRWs with monotonically decreasing variance—our primary focus in the present paper—is motivated by and closely related to two-dimensional directed polymers (see Section 4 for a detailed explanation of this relationship). On the other hand, the model of Gaussian Multiplicative Chaos (GMC) has long been viewed as a continuous analogue of BRWs, and numerous works connect these two models: for instance, many results from one often inspire analogous results for the other, and many approaches used in BRWs can be applied to those in GMC-type objects (see, for example [18] for a survey). In the present paper, we aim to proceed in the opposite direction: we use recent developments in GMC to further the study of BRWs.
Nonetheless, the complexity of the underlying partition function is a more intricate topic, as the classical martingale structure no longer holds and the correlation function is not sufficiently regular. Such problems have been avoided in most literature on BRWs and directed polymers. However, recent developments in GMC extend the theory to the non-martingale construction of GMC measures. In particular, Berestycki’s paper [4] provides a novel perspective that focuses on analyzing the ”thick points” that dominate the total mass. Motivated by this framework, we can not only provide an alternate proof of the classical result regarding phase transition of BRWs but also perform a similar construction in the BRWs literature and establish the universal behavior of the partition functions: in the sub-critical regime, the two partition functions (for time-homogeneous and time-inhomogeneous environments) converge to the same non-degenerate random variable. Our technical framework not only unifies the phase transition analysis of homogeneous and time-inhomogeneous BRWs but also provides a reusable tool for martingale-free random processes, enriching the cross research between BRWs and GMC.
These counterintuitive results in the present paper suggests that, in such settings, the fundamental asymptotic property of the partition function is determined solely by the underlying branching process and proves robust to time inhomogeneity in the weight increments. The core of the approaches used in the present paper is the observation that the total weight is dominated by the ”early generations” of the BRW—a key universal property unreported for time-inhomogeneous BRWs, distinct from that in [14].
Looking ahead, the results in the present paper suggest that, within the framework of GMC, the correlation function of time-inhomogeneous BRWs falls outside the scope of log-correlation, which in turn implies that the GMC framework can be extended to more general settings. Additionally, motivated by the fundamental role of BRWs in various branching processes and directed polymer models, our study raises a broader question: How does time-inhomogeneity impact phase transitions in these systems? The authors believe that the framework discussed in the present paper can be extended to some of these models.
@@ 1.1 Outline:
In Section 2, we introduce the model of BRWs (in time-homogeneous and time-inhomogeneous environments), the GMC and their connections, and state our main results. In Section 3, we will prove Theorem 2.2, and we can see that Theorem 2.1 serves as a direct consequence of the proof. Then we will prove the upper bound of $\bar{\beta}_{c}$, and together it shows that $\bar{\beta}_{c}=\beta_{c}$. In Section 3.1, we extend the universality to the critical parameter, showing the same polynomial decay rate of partition functions for time-inhomogeneous BRWs 3.7. Finally, we will discuss our results in context and raise some potential directions for further study in Section 4.
@@ 1.2 Acknowledgment:
The author wishes to thank Prof. Clément Cosco for proposing this interesting model, for insightful discussions that led to Theorem 2 and Theorem 3.6, and for valuable comments on the early stages of this work. The author is also very grateful to Jiaqi Liu for discussions about this research in the context of branching processes.
2 The Model of Branching Random Walks and Main Results
In the present paper, we consider a branching random walk on the real line $\mathbb{R}$. Initially, a particle sits at the origin. Its children form the first generation; their displacements from the origin correspond to a point process on the line (which could also be described by a Galton-Watson process, abbreviated as G-W process or G-W trees). These children have children of their own (who form the second generation) and behave—relative to their respective positions—like independent copies of the initial particle, and so on. We will focus on time-homogeneous and time-inhomogeneous Gaussian Branching Random Walks (BRWs) on a supercritical Galton-Watson tree $\mathbb{T}$ rooted at $o\in\mathbb{T}$ with mean offspring distribution equal to $\mathbb{E}[D_{1}]=d>1$, and the detailed definition of the model will be introduced below. In the rest of the paper, we will use the survival probability of the G-W tree:
@@ 2 The Model of Branching Random Walks and Main Results
Theorem 2.1.
@@
The classical proof of the previous result (see, for example, Chapter 5 of [17]) relies on a so-called ”spine method” (or ”spine decomposition”), an extensively used technique in the branching process literature, where one uses a change-of-measure argument and tilts the measure towards a particle chosen uniformly at random. Leveraging the martingale property, it turns out that under the tilted measure, the chosen particle dominates the total mass.
@@
The relationship between BRWs and GMC can be briefly described as follows (roughly speaking, we embed a binary BRWs into the unit interval, which yields so-called ”multiplicative cascades”, a preliminary version of GMC): First, let $\tilde{I}$ denote the infinite union of sub-intervals of $[0,1]$ defined by
$\tilde{I}=\bigcup_{m\geq 1}\tilde{I}_{m}=\bigcup_{m\geq 1}\bigcup_{i=0}^{2^{m}-1}\left[\frac{i}{2^{m}},\frac{i+1}{2^{m}}\right],$
where $\tilde{I}_{m}$ denotes the set of $m$-th level sub-intervals. We then assign a random variable $W_{I}$ to each interval $I\in\tilde{I}$: for any $I\in\tilde{I}_{k}$ (with $k\in\mathbb{N}\cup\{0\}$), set $W_{I}=e^{\beta W_{I}^{\prime}-\frac{\beta^{2}}{2}}$ for some $\beta>0$, where $W_{I}^{\prime}\sim\mathcal{N}(0,1)$ and the $W_{I}^{\prime}$ are independent across distinct $I\in\tilde{I}$.
For any point $x\in[0,1]$, define $w_{n}(x)=\prod_{I\in\tilde{I},\,x\in I}W_{I}$ (equivalently, $w_{n}(x)=\prod_{I\in\tilde{I}}W_{I}\cdot\mathbf{1}_{\{x\in I\}}$, where $\mathbf{1}_{\{\cdot\}}$ denotes the indicator function). We then introduce the random measure
$\mu_{n}:=\mu(n,\beta)=w_{n}\cdot\lambda,$
where $\lambda$ is the Lebesgue measure on $[0,1]$. For each $n\geq 1$, $\mu_{n}$ defines a random measure (not necessarily a probability measure) on $[0,1]$, and by independence of $W_{I}$ we can see that $\mathbb{E}[\mu_{n}([0,1])]=\prod_{I\in\tilde{I}_{n}}\mathbb{E}[W_{I}]=1$ for all $n$. This construction can be extended to GW trees via analogous ideas.
@@
$\mu(dx)=e^{\gamma X(x)-\frac{1}{2}\gamma^{2}\mathbb{E}[X(x)^{2}]}\,\sigma(dx).$ (1)
@@
  • $h$ is a centered Gaussian generalized random field with covariance kernel $K(x,y)=-\log|x-y|+g(x,y)$, where $g(x,y)$ is continuous on $\bar{D}\times\bar{D}$;
@@
In Berestycki’s paper [4] provides an elementary proof for the convergence of approximating measures $\mu_{\epsilon}$ constructed via the convolution of $h(x)$ with certain regularizing kernels $\theta_{\epsilon}$ to a non-trivial limiting measure $\mu$ in the sub-critical regime (specifically, when the parameter satisfies $\gamma<\sqrt{2d}$) and further demonstrates that this limiting measure $\mu$ is independent of the choice of regularizing kernels. The GMC framework with thick point analysis, i.e. the points $x\in D$ satisfying:
@@
which was shown in [4] that they dominates the total GMC measure, and measure tilting is a more suitable technical tool for phase transition analysis for BRWs in time-inhomogeneous environments. In time-inhomogeneous environment, the weight increment of any branch $\mathbf{|x|}=n$ exhibits a macroscopic variation: let $f:[0,1]\to[0,1]$ be continuous at 0, monotone decreasing function satisfying $f(0)=1$ and $f(1)=0$. For each generation in G-W tree $h\in\{0,...,n\}$, the weight increment at step $h$ follows the distribution of $\sqrt{f(\frac{h}{n})}w(x_{h})$. In other words, the Hamiltonian of a branch $\mathbf{x}=\{x_{1},...,x_{n}\}$, denoted by $\bar{H}_{n}(\mathbf{x})$, is given by: $\bar{H}_{n}(x)=\sum_{i=1}^{n}\sqrt{f(\frac{i}{n})}w(x_{i})$, so that $\bar{H}_{n}(\mathbf{x})\sim\mathcal{N}(0,\sum_{i=1}^{n}f(\frac{i}{n}))$.
Analogous to the normalized partition function for the BRW with homogeneous weight increments, we define the normalized partition function for the time-inhomogeneous BRW of length $n$, denote $\bar{W}_{n}^{\beta}$, as follows:
@@
In the present paper, we will establish the following stronger result: when $\beta<\beta_{c}$, the difference $\bar{W}_{n}^{\beta}-W_{n}^{\beta}\to 0$ in the $L^{1}$-norm, which in turn implies convergence to 0 in probability. Notably, $\bar{W}_{\infty}^{\beta}$ follows the same law as $W_{\infty}^{\beta}$-the a.s. limit of partition function in homogeneous weights. As a direct consequence, we can see that $\bar{W}_{\infty}^{\beta}$ is almost surely non-negative when $\beta<\sqrt{2\log d}$:
@@
Remark: Indeed, from the proof in Section 3 we can see that the restriction of $f$ in Theorem 2.2 could be relaxed: the proof will hold as long as $f:[0,1]\to[0,1]$ with $f(0)=1$ and being continuous at 0 (so it serves as a generalization of the condition of the decreasing variance in [14]), and the universality could be easily extended to any finite dimensional marginals.
Finally, the exact asymptotic behavior at criticality was established in the influential work of Hu and Shi,who proved that at the critical parameter $\beta_{c}$, $W_{n}^{\beta}$ possess the asymptotic behavior:
@@
(Asymptotics of partition function at criticality) At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.

### 27 Mar 2026 v2

@@ Abstract
We establish the phase transition and universality for the partition function of time-inhomogeneous branching random walks (BRWs) with decreasing variance increment, a model related to two-dimensional directed polymers. By modifying Berestycki’s GMC framework and adapting it to discrete BRWs, we prove that the critical value of time-inhomogeneous BRWs coincides with that of time-homogeneous ones, and the partition functions converge in $L^{1}$ in the subcritical regime. We also extend the universality at the critical parameter, showing the same decay rate of partition functions. Our approach reveals the potential for such framework beyond GMC, which provides a new technical path for other models, such as BRWs and Continuous Random Energy Model (CREM). Finally, we raise some open problems regarding the further implications of our results in other related models discussed in the present paper.
@@ 1 Introduction
Branching random walks (BRWs) are a natural model in probability theory, motivated by their biological applications, that describe the evolution of a population of particles with spatial motion. Much is known about the phase transition of their partition function (the so-called Biggins martingale); see [24] for precise statements and proofs. More recently, their time-inhomogeneous counterpart, first introduced in [19], initiates the study of the asymptotic behaviors of BRWs with increments that exhibit a macroscopic change in variance, revealing differences in extreme value statistics. The series of works on this model raises a more fundamental question: How do environmental fluctuations impact the asymptotic behavior of BRWs? In the present paper, we attempt to answer this question from a qualitative perspective by studying its partition function.
@@
The partition function of time-inhomogeneous BRWs with monotonically decreasing variance—our primary focus in the present paper—is motivated by and closely related to two-dimensional directed polymers. On the other hand, the model of Gaussian Multiplicative Chaos (GMC), originaed in [17] has long been viewed as a continuous analogue of BRWs, and numerous works connect these two models: for instance, many results from one often inspire analogous results for the other, and many approaches used in BRWs can be applied to those in GMC-type objects (see, for example [22] and [25] for some surveys of the connections between BRWs and two-dimensional Gaussian Free Field, a special case of GMC). In the present paper, we aim to proceed in the opposite direction: we use the framework in recent developments in GMC to further the study of BRWs.
We reveal the complexity of the partition function in time-inhomogeneous environments, a more intricate topic since the classical martingale structure fails and the correlation function lacks sufficient regularity—issues largely avoided in existing literature on BRWs and directed polymers. Fortunately, recent advances in GMC extend the theory to non-martingale constructions of GMC measures. In particular, Berestycki [4] offers a novel perspective by analyzing the “thick points” (roughly, points with a prescribed growth rate in the approximation) that dominate the total mass.
Indeed, this framework has been applied to related objects in the GMC literature (with significant technical differences across distinct settings), including convergence of critical GMC under convolution approximation in [21] and derivatives of subcritical GMC in [16]. Our results demonstrate the broad applicability of the framework from [4], which extends to models beyond GMC.
Motivated by this framework, we give an alternative proof of the classical phase transition result for BRWs, carry out an analogous construction in the BRW setting, and establish universal behavior of the partition functions: in the subcritical regime, the time-homogeneous and time-inhomogeneous partition functions converge to the same non-degenerate random variable. These results imply that, when the system behaves regularly in early-time evolution, the fundamental asymptotic property of the partition function is determined solely by the underlying branching process and is robust to time inhomogeneity in weight increments. The core of our approach is that fluctuations of the partition function are dominated by the early generations of the BRW.
Indeed, such a finding has been verified in a closely related model, the continuous random energy model (CREM), first introduced in [6], where a phase transition of free energy occurs with the critical parameter determined by the initial correlation. In the present paper, we further investigate this phase transition regarding the non-degeneracy of the limit, which serves as a direct consequence of the framework discussed herein (and will be elaborated on in Section 3.2.2), where we provide a characterization of the non-degeneracy of the limit in the weak-disorder regime under certain conditions on the correlation structure.
Looking ahead, results in this paper suggest that within the GMC framework, the correlation function of time-inhomogeneous BRWs is not log-correlated, implying that the GMC framework can be extended to more general settings. Additionally, our results on the time-inhomogeneity of increments align with a fundamental question in [7] regarding how environmental fluctuations affect phase transitions (or other qualitative behavior) in disordered systems; see [8] for a recent survey on 2D stochastic heat flow (a continuous analogue of two-dimensional directed polymers) related to this question.
Heuristically, disordered systems are divided into two categories: disorder relevant and disorder irrelevant. Disorder relevance means arbitrarily small disorder (or random environment) changes the model’s large-scale qualitative behavior (e.g., critical exponent changes) and leads to a different scaling limit; disorder irrelevance means such changes occur only for sufficiently strong disorder. At the boundary (typically critical dimension), disorder perturbation has a more subtle effect, being either marginally relevant or irrelevant. This paper also shows that the partition function has the same asymptotic behavior as its time-homogeneous counterpart at the critical parameter.
@@ 1.1 Outline:
In Section 2, we introduce the model of BRWs (in time-homogeneous and time-inhomogeneous environments), state our main results, as well as the potential connections to related models, such as GMC/two-dimensional directed polymers and CREM. In Section 3.1, we will prove Theorem 2.2. Then we will prove the upper bound of the critical parameter $\bar{\beta}_{c}$ in Section 10, and together these results show that $\bar{\beta}_{c}=\beta_{c}$. Then we will discuss the application of our main result, from which we can see that Theorem 2.1 serves as a direct consequence of the proof, and we can also extend the idea to the CREM model, as discussed in Proposition 3.2.2. In Section 3.3, we extend the universality to the critical parameter, showing the same polynomial decay rate of partition functions for time-inhomogeneous BRWs (Theorem 2.3). Finally, we will discuss our results in context and raise some other potential directions for further study in Section 4.
@@ 1.2 Acknowledgment:
The author wishes to thank Clément Cosco for proposing this interesting model, as well as discussions and comments on the early stages of this work. The author is also very grateful to Jiaqi Liu and Bastein Mallein for discussions about these results in the broader context of modern research.
2 Descriptions of Models and Main Results
2.1 The Model of Branching Random Walks
In the present paper, we consider a branching random walk on the real line $\mathbb{R}$. Initially, a particle sits at the origin. Its children form the first generation; their displacements from the origin correspond to a point process on the line (which could also be described by a Galton-Watson process, abbreviated as G-W process or G-W trees). These children have children of their own (who form the second generation) and behave—relative to their respective positions—like independent copies of the initial particle, and so on. We will focus on time-homogeneous and time-inhomogeneous Gaussian Branching Random Walks (BRWs) on a supercritical Galton-Watson tree $\mathbb{T}$ rooted at $o\in\mathbb{T}$ with mean offspring distribution equal to $\mathbb{E}[D_{1}]=d>1$ (we denote $D_{n}$ to be the particles at generation $n$), and the detailed definition of the model will be introduced below. In the rest of the paper, we will use the survival probability of the G-W tree:
@@ 2.1 The Model of Branching Random Walks
Theorem 2.1 (Biggins Martingale Convergence Theorem).
@@
The classical proof of the previous result (see, for example, Chapter 5 of [24]) relies on a so-called “spine method” (or “spine decomposition”), an extensively used technique in the branching process literature, where one uses a change-of-measure argument and tilts the measure towards a particle chosen uniformly at random. Leveraging the martingale property, it turns out that under the tilted measure, the chosen particle dominates the total mass.
2.2 Gaussian multiplicative Chaos
@@
The relationship between BRWs and GMC can be briefly described as follows (roughly speaking, we embed a binary BRWs into the unit interval, which yields so-called “multiplicative cascades”, a preliminary version of GMC): First, let $\tilde{I}$ denote the infinite union of sub-intervals of $[0,1]$ defined by
$\tilde{I}=\bigcup_{m\geq 1}\tilde{I}_{m},\text{ where }\tilde{I}_{m}:=\bigcup_{i=0}^{2^{m}-1}\left[\frac{i}{2^{m}},\frac{i+1}{2^{m}}\right],$
Then for any $I\in\tilde{I}_{k}$ (with $k\in\mathbb{N}\cup\{0\}$), set $W_{I}=e^{\beta W_{I}^{\prime}-\frac{\beta^{2}}{2}}$ for some $\beta>0$, where $W_{I}^{\prime}\sim\mathcal{N}(0,1)$ and the $W_{I}^{\prime}$ are independent across distinct $I\in\tilde{I}$. For any point $x\in[0,1]$, define $w_{n}(x)=\prod_{I\in\tilde{I}_{n},\,x\in I}W_{I}$. We then introduce the random measure $\mu_{n}:=\mu(n,\beta)$ satisfying: $\frac{\text{d}\mu_{n}}{\text{d}\lambda}=w_{n}$, where $\lambda$ is the Lebesgue measure on $[0,1]$. For each $n\geq 1$, $\mu_{n}$ defines a random measure on $[0,1]$, and by independence of $W_{I}$ we can see that $\mathbb{E}[\mu_{n}([0,1])]=\prod_{I\in\tilde{I}_{n}}\mathbb{E}[W_{I}]=1$ for all $n$. This construction can be extended to GW trees via analogous ideas (according to each sampling of G-W tree), which will be used in the proof of Theorem 2.3.
@@
$\mu(dx)=e^{\gamma h(x)-\frac{1}{2}\gamma^{2}\mathbb{E}[h(x)^{2}]}\,\sigma(dx).$ (1)
@@
  • $h$ is a centered Gaussian generalized random field with covariance kernel $K(x,y)=\log|x-y|^{-1}+g(x,y)$, where $g(x,y)$ is continuous on $\bar{D}\times\bar{D}$;
@@
In Berestycki’s paper [4], the author provides an elementary proof for the convergence of approximating measures $\mu_{\epsilon}$ constructed via the convolution of $h(x)$ with certain regularizing kernels $\theta_{\epsilon}$ to a non-trivial limiting measure $\mu$ in the sub-critical regime (specifically, when the parameter satisfies $\gamma<\sqrt{2d}$) and further demonstrates that this limiting measure $\mu$ is independent of the choice of regularizing kernels. The GMC framework with thick point analysis, i.e. the points $x\in D$ satisfying:
@@
which was shown in [4] Lemma 3.4 that they dominates the total GMC measure, and measure tilting argument used in that paper will serve a more suitable technical tool for phase transition analysis for BRWs in time-inhomogeneous environments, which will be the main focus in the present paper. The later model was initially discussed in Fang and Zeitouni’s paper [19] and drew attention from researchers in multiple related areas. In particular, recent study in [11] suggests that there is an analogy between two-dimensional directed polymers and BRWs with monotone decreasing variance, which serves as one of the primary motivation for the present paper.
2.3 Two-dimensional Directed Polymers and Time-inhomogeneous BRWs
Directed polymers in random environment describe the behavior of a long directed chain of monomers in presence of random impurities, see [10] for a comprehensive introduction to this area. In the setting related to our discussions, the trajectory of the polymer is given by a nearest-neighbor path $(S_{n})_{n\in\mathbb{N}}$ on two-dimensional lattice, while the impurities (also called the environment) are given by i.i.d. centered random variables $\omega(n,x)$, $n\in\mathbb{N},x\in\mathbb{Z}^{d}$ of variance one satisfying the free energy $\lambda(\beta)=\log\mathbb{E}[e^{\beta\omega(n,x)}]\in(0,\infty)$ for all $\beta>0$, and define $W_{N}(x,\beta)=\mathbb{E}_{x}\left[e^{\sum_{n=1}^{N}\left(\beta\omega(n,S_{n})-\lambda(\beta)\right)}\right]$ to be the partition function for paths start at $x$. In the seminal paper [7], Caravenna, Sun and Zygouras established the convergence of the spatial behavior of two-dimensional (re-scaled) partition functions to a Gaussian random variable (see [7], Theorem 2.5 for a precise statement).
In [11], the authors showed that when the environment are i.i.d. standard Gaussian (i.e. having law $\mathcal{N}(0,1)$), the normalized partition function can be heuristically viewed as a branching random walk with decreasing variance (where each branch corresponds to a point $x\in\mathbb{Z}^{d}$), satisfy that $X_{k}(x)$ and $X_{k}(y)$ are correlated for small |xy||x-y| ($x,y\in\mathbb{Z}^{2}$), and decorrelate as |xy||x-y| grows large. Consequently, the asymptotic behavior of the partition function should be governed by a branching random walk model with decreasing variance discussed in the present paper.
In time-inhomogeneous environment, the weight increment of any branch $\mathbf{|x|}=n$ exhibits a macroscopic variation: let $f:[0,1]\to[0,1]$ be continuous at 0, monotone decreasing function satisfying $f(0)=1$ and $f(1)=0$. For each generation in G-W tree $h\in\{0,...,n\}$, the weight increment at step $h$ follows the distribution of $\sqrt{f(\frac{h}{n})}w(x_{h})$. In other words, the Hamiltonian of a branch $\mathbf{x}=\{x_{1},...,x_{n}\}$, denoted by $\bar{H}_{n}(\mathbf{x})$, is given by: $\bar{H}_{n}(x)=\sum_{i=1}^{n}\sqrt{f(\frac{i}{n})}w(x_{i})$, so that $\bar{H}_{n}(\mathbf{x})\sim\mathcal{N}(0,\sum_{i=1}^{n}f(\frac{i}{n}))$. Analogous to the normalized partition function for the BRW with homogeneous weight increments, we define the normalized partition function for the time-inhomogeneous BRW of length $n$, denote $\bar{W}_{n}^{\beta}$, as follows:
@@ 2.3 Two-dimensional Directed Polymers and Time-inhomogeneous BRWs
In the present paper, we will establish the following stronger result: when $\beta<\beta_{c}$, the difference $\bar{W}_{n}^{\beta}-W_{n}^{\beta}\to 0$ in the $L^{1}$-norm, which in turn implies convergence to 0 in probability. Notably, the limit of $\bar{W}_{n}^{\beta}$ ,denoted by $\bar{W}_{\infty}^{\beta}$, follows the same law as $W_{\infty}^{\beta}$-the a.s. limit of partition function in time-homogeneous weights. As a direct consequence, we can see that $\bar{W}_{\infty}^{\beta}$ is almost surely non-negative when $\beta<\sqrt{2\log d}$:
@@
Finally, the exact asymptotic behavior at criticality was established in the influential work of Hu and Shi [15],who proved that at the critical parameter $\beta_{c}$, $W_{n}^{\beta}$ possess the asymptotic behavior:
@@
[Asymptotics of partition function at criticality] At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.
@@
2.4 Continuous Random Energy Model
Indeed, while the property of $f$ in the statement of Theorem 2.2 was our primary motivation from two-dimensional directed polymers, from the proof in Section 3.1 we can see that the restriction on $f$ in Theorem 2.2 could be relaxed (i.e., Theorem 2.2 holds for a generalized class of $f$): the proof will hold as long as $f:[0,1]\to[0,1]$ with $f(0)=1$, is continuous at $x=0$.
Motivated by the mild conditions on $f$, our result could also be extended to another closely related model: the continuous random energy model (CREM), which was introduced by Bovier and Kurkova in the paper [6] in 2004, which serves as another toy model of disordered systems exhibiting an infinitely hierarchical correlation structure; see also [14] and [18] for more recent studies of this model. The model is defined as a generalized version of binary time-homogeneous BRWs: denote by $\mathbb{T}_{n}$ the binary tree with depth $n$, and given $\mathbf{x},\mathbf{y}\in\mathbb{T}_{n}$, $C(\mathbf{x},\mathbf{y})$ is the depth of their most recent common ancestor (see the precise definition in the proof of Theorem 3). The CREM is a centered Gaussian process indexed by the binary tree $\mathbb{T}_{n}$ of depth $N$ (or equivalently, on the hypercube $\{0,1\}^{N}$ in other literature, and it serves as a toy model of spin glass on $\{0,1\}^{N}$ in this case) with covariance function:
$\text{Cov}\left(\mathbf{x},\mathbf{y}\right)=n\cdot A\left(\frac{C(\mathbf{x},\mathbf{y})}{n}\right),\quad\forall|\mathbf{x}|,|\mathbf{y}|=n.$
where $A:[0,1]\to[0,1]$ is a non-decreasing function satisfying $A(0)=0$ and $A(1)=1$, and the right derivative of $A$ at 0, denoted by $\hat{A}^{\prime}(0)$, is finite (i.e., $\hat{A}^{\prime}(0)<\infty$).
To study the CREM, one of the key quantities is the normalized partition function defined as $Z_{n}^{\beta}$ (the quantity $\beta_{c}$ is sometimes referred to as the static critical inverse temperature of the CREM):
$Z_{n}^{\beta}:=\frac{\sum_{|\mathbf{x}|=n}e^{\beta H_{n}(\mathbf{x})}}{\mathbb{E}[\sum_{|\mathbf{x}|=n}e^{\beta H_{n}(\mathbf{x})}]}$ (2)
From the expression of the covariance function, we can see that it possesses a similar construction to time-inhomogeneous BRWs on binary trees. Indeed, one may view $A(x)$ as $\int_{0}^{x}f(t)dt$ in the time-inhomogeneous BRWs for any $x\in[0,1]$, which serves as the limit of the discrete sum of $f(\frac{i}{n})$ (whereas the case $A(x)=x$ corresponds to time-homogeneous BRWs). In the same paper, Bovier and Kurkova showed that there exists a critical parameter $\beta_{c}=\frac{\sqrt{2\log 2}}{\sqrt{\hat{A}^{\prime}(0)}}$ such that the following phase transition occurs: heuristically speaking, for all $\beta<\beta_{c}$, the main contribution to the partition function $Z_{n}^{\beta}$ comes from an exponential number of the particles, and for all $\beta>\beta_{c}$, the maximum starts to contribute significantly to the partition function.
However, to our knowledge, it has not yet been proven in the literature that $\beta_{c}$ serves as the critical parameter for the non-degeneracy of the limit of $Z_{n}^{\beta}$. Under the framework of the present paper, we can see that although the exact approximation terms of $Z_{n}^{\beta}$ are different from those of $\bar{W}_{n}^{\beta}$, the similarity in Gaussian increments and the covariance structure allows us to prove the convergence of the partition function under certain conditions on $A$:
Proposition 2.4 (Convergence of CREM Partition function).
Suppose for any $x\in[0,1]$, $A^{\prime}(x)$ exists almost everywhere, satisfies $\sup_{x\in[0,1]}A^{\prime}(x)\leq A^{\prime}(0)$, and is continuous at 0 (i.e., $A\in C^{0}[0,1]\cap C^{1}(0)$). Then when $\beta<\beta_{c}=\frac{\sqrt{2\log 2}}{\sqrt{A^{\prime}(0)}}$, we have $Z_{n}^{\beta}$ converges to $Z_{\infty}^{\beta}$ in $L^{1}$ and $Z_{\infty}^{\beta}$ has the same law as $W_{\infty}^{\beta}$.
Many other interesting results have been established for the CREM and its partition function, and we believe that it would be interesting to compare those results with the BRWs literature (including the results in the present paper).